OpenStudy (anonymous):

9x-4+3(x-11)=23 find x step by step

OpenStudy (merchandize):

9x-4+3x-33=23 12x=23+33+4 12x=60 x=60/12 x=5

OpenStudy (anonymous):

9x-4+3(x-11)=23 9x+3x-33=23+4 9x+3x=27+33 12x=60 12x/12=60/12 x=5